A 25.0 milliliters sample of HNO3 (aq) is neutralized by32.1 milliliters of 0.150 M KOH (aq). What is the concentration of the acid? Show a numerical setup

1
Expert's answer
2022-04-17T11:04:28-0400

HNO3 + KOH = KNO3 + H2O

Cm = nV

n(KOH) = 0.150M / 0.0321L = 4.67 mol

n(KOH) = n(HNO3) = 4.67 mol

C(HNO3) = 4.67 mol * 0.025 L = 0.117M